# Specialization after Generalization: Towards Understanding Test-Time Training in Foundation Models

### v2

@@ E.2 SAE on MNIST data
The procedure of searching for the optimal mask $m$ stays the same as per Section 3, however, we note that in this setup the penalty $\lambda$ is set to $10^{-3}$. For this experiment the average resulting sparsity of the mask $m$ is $\mathbb{E}_{x^{*}}[\|m\|_{0}]\approx 24.5.$
For a dense base model (i.e., a counterpart of CLIP embeddings in case of MNIST), we train a variant of LeNet (LeCun et al., 1998a) with a scale of $0.5$. The resulting feature dimension before the final linear classification layer is equal to $50$. This model is clearly underparameterized with test accuracy $94.51\%$.
For the base CNN TTT model, we perform 100 full-batch steps of Adam with learning rate of $10^{-1}$ (increasing budget does not affect the performance). For TTT with the adaptive mask in concept space, we do 200 full-batch steps of Adam with learning rate $5\cdot 10^{-2}$. The neighborhood size for TTT is set to $n=100$.
The results in terms of accuracy are as follows:
@@
We use the same notation for active sets for concept vectors as in Section˜3. However, for better clarity, we will recap the notation. For each sparse vector $\Phi(x)\in\mathbb{R}^{d_{1}}$ we define its active set as the set of vector components that are non-zero, i.e.,
@@
In these terms, the “intersection” analysis reveals the following properties of the adaptive mask $m$, test active set $m_{x^{\star}}$ and neighbors’ active sets $m_{i}$:
@@
In particular, it might be tempting to draw the following simplifying conclusion: why not just pick
@@
i.e., let mask $m$ only select indices that appear both in the test point and its neighbors? The flaw of this approach becomes evident for more complex data (e.g., ImageNet in Section˜3), as slack components of the mask are necessary to select “non-spurious” test components.
E.3 SAE on ImageNet CLIP embeddings
@@
Obtaining a sparse autoencoder with meaningful features (with mild amount of non-active neurons) is a task with multiple caveats. We use a few common techniques that make the training procedure more successful, which we describe below.
We use learning rate warm-up to ensure that in the beginning of training the concept space is explored properly, i.e., neurons do not deactivate and get stuck at a suboptimal configuration. We warm up the learning rate linearly to the value $3\cdot 10^{-4}$ for $T_{0}=5000$ steps. After this, we employ a typical cosine decay with a horizon of $T=10^{5}$. In particular, let $i$ be the current step; then at each step the initial learning (in this case $3\cdot 10^{-4}$) is multiplied by the value of $\lambda_{i}$, which is computed as follows:
$\tilde{\lambda}_{i}=(i-T_{0})/(T-T_{0})$
[…]
@@
$\lambda_{i}=i/K$
$k_{i}=k_{0}-(k_{0}-k)\cdot\lambda_{i}.$
@@
We now describe the implementation of the ghost gradients Gao et al. (2025) used in our ImageNet run. In a nutshell, ghost loss ensures that features that are not activated in the top-$k$ are still getting learning signal during the training. This is very important as non-convexity of the optimization landscape often leads to a suboptimal configuration for which considerable amount of units is inactive. Ghost gradients method aims at “shaking” these units up to make them active again. Thus, we first need to define which units are considered inactive during the training. Let $f_{i}\in\mathbb{R}^{d_{1}}$ be the vector of frequencies at iteration $i$ for activity of each unit, i.e.,
@@
We add ghost loss (23) to the initial reconstruction objective with weight of $10^{6}$.
@@
E.4 Scaling experiments on MNIST
In the scaling experiments reported in Section˜4, we trained multiple LeNet (LeCun et al., 1998a) convolutional neural networks (CNNs) at different model scales. The architecture comprises two convolutional layers, each with ReLU activation and $2\times 2$ max-pooling, followed by a fully connected classification head. The reference models of varying sizes were trained with the hyperparameters listed in Table˜3, obtained by tuning on the validation set. Optimization was performed with Adam (Kingma & Ba, 2015).
@@ E.4 Scaling experiments on MNIST
Hyperparameter 280 600 1268 1976 2985 4430 6386 11770 24294 40818
Learning Rate 6e-3 2e-3 1e-3 8e-4 6e-4 6e-4 2e-3 2e-3 2e-3 6e-4
Batch Size 500 200 400 600 300 300 400 300 300 100
Epochs 50 50 100 100 100 100 100 100 100 50
@@
For the model scaling plot in Figure˜4, we trained each reference model across five random seeds and applied both TTT and majority voting. For TTT, all model parameters were frozen except for the final linear layer, which was fine-tuned from its pre-trained initialization using the hyperparameters in Table˜4. Majority voting was performed with neighborhood sizes specified in Table˜5. As the experiments were repeated over five seeds, we used 200 bootstrap iterations per seed to compute confidence intervals.
@@
Hyperparameter 280 600 1268 1976 2985 4430 6386 11770 24294 40818
Learning rate 0.05 5e-3 0.01 0.01 5e-3 5e-3 1e-3 5e-3 1e-3 1e-3
Epochs 50 200 200 200 200 200 200 200 200 200
Number of neighbors 200 50 10 200 200 200 50 100 50 200
@@ E.5 Connection to MoEs
Given an underparameterized global model, a natural alternative to specializing to each test-time task as in TTT is to instead specialize individual “expert” models to subsets of tasks, routing test-time tasks to few of these experts. Such mixture of experts (MoEs; Shazeer et al., 2017; Fedus et al., 2022; Bertolissi et al., 2025) have been shown to be an effective architecture for foundation models (e.g., Dai et al., 2024). To see whether our findings extend to MoEs, we train multiple experts (each being a different linear head) based on the MoE architecture of Bertolissi et al. (2025), and evaluate accuracy as we scale the number of experts. We find that a larger number of experts increases the capacity of the model and improves accuracy, highlighting that MoEs are a promising approach to specialization without increasing inference cost.
We used a pre-trained CNN to extract last-layer embeddings from MNIST images. Following Bertolissi et al. (2025), for a given number of experts, each expert was associated with a cluster centroid obtained by partitioning the training set into clusters via $k$-means. Expert training was performed by fine-tuning the pre-trained linear head on the nearest neighbors of the assigned centroid in embedding space. The fine-tuning hyperparameters are provided in Table˜6. At test time, inputs were first mapped through the CNN encoder, after which the single closest cluster centroid was selected based on $L_{2}$-distance. The corresponding expert head then produced the final prediction.
@@
E.6 Scaling experiments on ImageNet
@@
Building upon the baseline linear classifier, we apply a test-time training (TTT) procedure to adapt the model at inference-time. Specifically, we fine-tune this linear head for each test sample by minimizing the cross-entropy loss over the set of $k=600$ nearest neighbors retrieved in the original CLIP embedding space. Optimization proceeds for 50 epochs using the Adam optimizer with a learning rate of 0.02 and a batch size equal to the number of neighbors.
@@
To assess the impact of embedding dimensionality, we trained two-layer multi-layer perceptrons (MLPs) as classification heads on top of the CLIP embeddings. The first hidden layer had variable size, ranging from 250 to 4000 neurons, followed by a ReLU activation. The second layer maps the hidden representation to the 1000 ImageNet classes. Dropout is applied before the final layer to mitigate overfitting. Training was conducted with the Adam optimizer, using the hyperparameters specified in Table˜7.
@@
After training, we freeze the first hidden layer and project the original 512-dimensional embeddings into this hidden space. The resulting hidden representations, taken after the ReLU activation, serve as our scaled embeddings for test-time training. Specifically, we fine-tune the pre-trained linear MLP head on the set of $k$ nearest neighbors by minimizing the cross-entropy loss using Adam. This setup allows us to systematically examine how performance and representation quality vary with embedding dimensionality. We used a set of universal hyperparameters that are nearly optimal across all model sizes, as reported in Table˜8. Loss optimization was performed in a full-batch setting, with the batch size equal to the number of neighbors.
@@
As an alternative baseline, we leverage the learned feature spaces of the two-layer MLPs and apply a simple majority-voting protocol based on nearest neighbors. Specifically, we first map the original 512-dimensional CLIP embeddings into the MLP’s hidden feature space and then identify its $k=10$ nearest neighbors for any given test sample. The predicted class is assigned as the most frequent (majority) class label among these neighbors. This approach parallels the neighbor selection used in test-time training (TTT) but replaces fine-tuning with a straightforward plurality vote. Majority voting thus serves as a simple, non-parametric baseline to assess the quality of the scaled embeddings, providing insight into the clustering and class separability properties of the learned feature space.
E.7 Scaling experiments for language modeling
We use the open-source implementation [^9] https://github.com/socialfoundations/tttlm of Hardt & Sun (2024) and evaluate the Qwen2.5 family of base models (Qwen et al., 2025). We summarize hyperparameters in Table˜9. Sequences in the neighborhood that exceed the maximum sequence length are split into chunks of the maximum length. This means that a single neighbor can result in multiple gradient steps during TTT.

### v3

@@ F.2 SAE on MNIST data
The procedure of searching for the optimal mask $m$ stays the same as per Section 3, however, we note that in this setup the penalty $\lambda$ is set to $10^{-3}$. For this experiment, the average resulting sparsity of the mask $m$ is $\mathbb{E}_{x^{*}}[\|m\|_{0}]\approx 24.5.$
For a dense base model (i.e., a counterpart of CLIP embeddings in the case of ImageNet), we train a variant of LeNet (LeCun et al., 1998a) with a scale of $0.5$. The resulting feature dimension before the final linear classification layer is equal to $50$. This model is clearly underparameterized with test accuracy $94.51\,\%$.
For the base CNN TTT model, we perform 100 full-batch steps of Adam with learning rate of $10^{-1}$ (increasing budget does not affect the performance). For TTT with the adaptive mask in the concept space, we do 200 full-batch steps of Adam with learning rate $5\cdot 10^{-2}$. The neighborhood size for TTT is set to $n=100$.
The resulting accuracies are:
@@
We use the same notation for active sets for concept vectors as in Section˜3. However, for better clarity, we recap the notation. For each sparse vector $\Phi(x)\in\mathbb{R}^{d_{1}}$ we define its active set as the set of vector components that are non-zero, i.e.,
@@
In this context, the “intersection” analysis reveals the following properties of the adaptive mask $m$, active set of the test point $m_{x^{\star}}$ and neighbors’ active sets $m_{i}$:
@@
In particular, one might be tempted to draw the following simplifying conclusion:
@@
that is, to define the mask $m$ so that it only selects indices appearing in both the test point and its neighbors. However, this approach fails on more complex datasets (e.g., ImageNet, see Section˜3), because additional slack components in the mask are necessary to capture “non-spurious” test features.
F.3 SAE on ImageNet CLIP embeddings
@@
Obtaining a sparse autoencoder with meaningful features (with mild amount of non-active neurons) is a task with multiple caveats. We employ several common techniques to improve the training procedure, which we describe below.
We use a learning rate warm-up to ensure that the concept space is properly explored at the start of training, preventing neurons from deactivating and becoming trapped in suboptimal configurations. We warm up the learning rate linearly to the value $3\cdot 10^{-4}$ for $T_{0}=5000$ steps. After this, we employ a typical cosine decay with a horizon of $T=10^{5}$. In particular, let $i$ be the current step, then at each step the initial learning (in this case $3\cdot 10^{-4}$) is multiplied by the value of $\lambda_{i}$, which is computed as follows:
$\lambda_{i}:=0.5\cdot\left(1+\cos\left(\pi\cdot\tilde{\lambda}\right)\right)\qquad\text{with}\qquad\tilde{\lambda}_{i}:=(i-T_{0})/(T-T_{0}).$
@@
$k_{i}:=k_{0}-(k_{0}-k)\cdot\gamma_{i}\qquad\text{with}\qquad\gamma_{i}:=i/K.$
@@
We now describe the implementation of the ghost gradients (Gao et al., 2025) used in our ImageNet run. In a nutshell, ghost loss ensures that features that are not activated in the top-$k$ are still getting learning signal during the training. This is very important as non-convexity of the optimization landscape often leads to a suboptimal configuration for which considerable amount of units is inactive. The ghost gradient method aims at “shaking” these units up to make them active again. Thus, we first need to define which units are considered inactive during the training. Let $f_{i}\in\mathbb{R}^{d_{1}}$ denote the vector of unit activation frequencies after iteration $i$, where
@@
We add the ghost loss (22) to the initial reconstruction objective with weight of $10^{6}$.
@@
F.4 Scaling experiments on MNIST
In the scaling experiments reported in Section˜4, we trained multiple LeNet (LeCun et al., 1998a) convolutional neural networks (CNNs) at different model scales. The architecture comprises two convolutional layers, each with ReLU activation and $2\times 2$ max-pooling, followed by a fully connected classification head. The reference models of varying sizes were trained with the hyperparameters listed in Table˜4. Optimization was performed with Adam (Kingma & Ba, 2015).
@@ F.4 Scaling experiments on MNIST
280 600 1268 1976 2985 4430 6386 11770 24294 40818 61342 85866
Learning rate 6e-3 2e-3 1e-3 8e-4 6e-4 6e-4 2e-3 2e-3 2e-3 6e-4 2e-3 2e-3
Batch Size 500 200 400 600 300 300 400 300 300 100 300 500
Epochs 50 50 100 100 100 100 100 100 100 50 100 100
@@
For the model scaling plot in Figure˜4, we trained each reference model across five random seeds and applied both TTT and majority voting. For TTT, all model parameters were frozen except for the final linear layer, which was fine-tuned from its pre-trained initialization using the hyperparameters in Table˜5. Majority voting was performed with neighborhood sizes specified in Table˜6. All hyperparameters were obtained from a hyperparameter optimization on the validation set. As the experiments were repeated over five seeds, we used 200 bootstrap iterations per seed to compute confidence intervals.
@@
280 600 1268 1976 2985 4430 6386 11770 24294 40818 61342 85866
Learning rate 0.05 5e-3 0.01 0.01 5e-3 5e-3 1e-3 5e-3 1e-3 1e-3 5e-3 1e-3
Epochs 50 200 200 200 200 200 200 200 200 200 200 100
Neighbors 200 50 10 200 200 200 50 100 50 200 300 100
@@ F.5 Connection to MoEs
Given an underparameterized global model, a natural alternative to specializing to each test-time task as in TTT is to instead specialize individual “expert” models to subsets of tasks, routing test-time tasks to few of these experts. Such mixture of experts (MoEs; Shazeer et al., 2017; Fedus et al., 2022; Bertolissi et al., 2025) have been shown to be an effective architecture for foundation models (e.g., Dai et al., 2024).
To see whether our findings extend to MoEs, we train multiple experts (each being a different linear head) based on the MoE architecture of Bertolissi et al. (2025), and evaluate accuracy as we scale the number of experts. We find that a larger number of experts increases the capacity of the model and improves accuracy, highlighting that MoEs are a promising approach to specialization without increasing inference cost.
We used a pre-trained CNN to extract last-layer embeddings from MNIST images. Following Bertolissi et al. (2025), for a given number of experts, each expert was associated with a cluster centroid obtained by partitioning the training set into clusters via $k$-means. Expert training was performed by fine-tuning the pre-trained linear head on the nearest neighbors of the assigned centroid in embedding space. The fine-tuning hyperparameters are provided in Table˜7, obtained by tuning on the validation set. At test time, inputs were first mapped through the CNN encoder, after which the single closest cluster centroid was selected based on $L_{2}$-distance. The corresponding expert head then produced the final prediction.
@@
F.6 Scaling experiments on ImageNet
@@
Building upon the baseline linear classifier, we apply a test-time training procedure to adapt the model at inference-time. Specifically, we fine-tune this linear head for each test sample by minimizing the cross-entropy loss over the set of $k=600$ nearest neighbors retrieved in the original CLIP embedding space. By optimizing the error rate on the validation set, we determined that using $k=600$ neighbors with a learning rate of 0.02, a batch size equal to the number of neighbors, and 50 training epochs, combined with the Adam optimizer, perform near-optimally across all model scales. Consequently, we adopted this universal set of hyperparameters for the ImageNet experiments.
@@
To assess the impact of embedding dimensionality, we trained two-layer multi-layer perceptrons (MLPs) as classification heads on top of the CLIP embeddings. The first hidden layer had variable size, ranging from 250 to 4000 neurons, followed by a ReLU activation. The second layer maps the hidden representation to the 1000 ImageNet classes. Dropout is applied before the final layer to mitigate overfitting. Training was conducted with the Adam optimizer, using the hyperparameters specified in Table˜8, which were selected via optimization on the validation set.
@@
After training, we freeze the first hidden layer and project the original 512-dimensional embeddings into this hidden space. The resulting hidden representations, taken after the ReLU activation, serve as our scaled embeddings for test-time training. Specifically, we fine-tune the pre-trained linear MLP head on the set of $k$ nearest neighbors by minimizing the cross-entropy loss using Adam. This setup allows us to systematically examine how performance and representation quality vary with embedding dimensionality. The hyperparameters used for this procedure were obtained via hyperparameter optimization on the validation set, and we found that this universal set is nearly optimal across all model sizes, as reported in Table˜9. Loss optimization was performed in a full-batch setting, with the batch size equal to the number of neighbors.
@@
As an alternative baseline, we leverage the learned feature spaces of the two-layer MLPs and apply a simple majority voting protocol based on nearest neighbors. Specifically, we first map the original 512-dimensional CLIP embeddings into the MLP’s hidden feature space and then identify its $k=10$ nearest neighbors for any given test sample. Our empirical analysis showed that this neighborhood size performs well across all model scales. The predicted class is assigned as the most frequent (majority) class label among these neighbors. This approach parallels the neighbor selection used in test-time training (TTT) but replaces fine-tuning with a straightforward plurality vote. Majority voting thus serves as a simple, non-parametric baseline to assess the quality of the scaled embeddings, providing insight into the clustering and class separability properties of the learned feature space.
F.7 Scaling experiments for language modeling
We use the open-source implementation [^10] https://github.com/socialfoundations/tttlm of Hardt & Sun (2024) and evaluate the Qwen2.5 family of base models (Qwen et al., 2025). We summarize hyperparameters in Table˜10. During test-time training, we perform gradient steps sequentially for each sequence in the neighborhood. Sequences in the neighborhood that exceed the maximum sequence length are split into chunks of the maximum length. This means that a single neighbor can result in multiple gradient steps during TTT.